OpenStudy (anonymous):

What are the solutions of 3x2 + 4x = –5? the quantity negative 4 plus or minus i square root of 66 all over 6 the quantity negative 2 plus or minus 2i square root of 66 all over 3 the quantity negative 2 plus or minus i square root of 11 all over 3 the quantity negative 4 plus or minus i square root of 11 all over 6

OpenStudy (mertsj):

Did you put it equal to 0 and plug into the quadratic formula?

OpenStudy (mertsj):

3x^2+4x+5=0

OpenStudy (callisto):

3x^2 + 4x = –5 3x^2 +4x +5 =0 Now, use the quadratic formula for the equation. \[x= \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\] Put a =3, b=4 and c=5 into the formula. Can you do it?

OpenStudy (anonymous):

is it the first option?

OpenStudy (anonymous):

the quantity negative 4 plus or minus i square root of 66 all over 6?

OpenStudy (callisto):

Not really... Just consider the determinant b^2 - 4ac For a =3, b=4 and c=5 , b^2 - 4ac = 4^2 - 4 (3) (5) = 16 - 60 = ?